OpenStudy (anonymous):

solve by using substitution y=2x-4 7x-5y=14

OpenStudy (anonymous):

You want to turn the double variable equations into equations with one variable. We can do this by understanding that the two equations given are related. In this case, the preliminary work of solving for one variable in terms of the other is done for us, so we can substitute the y in the second equation with "2x-4", which was determined by solving the first equation (they already did it for us). We would then get: 7x-5(2x-4)=14 7x-10x+20=14 -3x+20=14 -3x=-6 x=2 Now that we know what x is, we can solve for y by plugging that value into the first equation: y=2(2)-4 y=0 And there you have it. I there is anything you didn't understand feel free to ask! I will try my best to explain! >.<

OpenStudy (anonymous):

i understand it now thanks(: